OpenStudy (anonymous):

The positive sum of two perfect squares is 32. What is the largest positive value of the sum of the two perfect square?

OpenStudy (dumbcow):

a^2 + b^2 = 32 ?

OpenStudy (anonymous):

sorry the question is should be "The positive difference of two perfect squares is 32. What is the largest positive value of the sum of the two perfect square?"

OpenStudy (dumbcow):

a = b = 4

OpenStudy (anonymous):

sorry

OpenStudy (turingtest):

oh, calculus?

OpenStudy (anonymous):

no, i don't think so

OpenStudy (anonymous):

\[a^2=b^2-32\]\[a, b \in \mathbb{Z}^+\]

OpenStudy (anonymous):

\[7^2+9^2=130\]

OpenStudy (mathmate):

(x^2-y^2)=(x+y)(x-y)=32 factors of 32: 1,2,4,8,16,32 For x & y to be integers, x+y & x-y must be both even, so 9^2-7^2=32, 9^2+7^2=130 6^2-2^2=32, but 6^2+2^2=30 So 9 & 7 are the numbers

OpenStudy (anonymous):

thx

OpenStudy (mathmate):

you're welcome! :)